OpenStudy (anonymous):

What is the slope of the line between (3, -4) and (-2, 1)?

OpenStudy (the_fizicx99):

M = (y2-y1)/(x2-x1) Plug in your values, M = 1 -(-4)/-2-3

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 3}}\quad ,&{\color{blue}{ -4}})\quad &({\color{red}{ -2}}\quad ,&{\color{blue}{ 1}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}}\)